# Gravitational Lensing of Euler-Heisenberg Black Hole Surrounded by Perfect Fluid Dark Matter

### v3

@@ 4.4 Black Hole Shadow
Figure 6: The black hole shadow of Euler-Heisenberg black hole surrounded by PFDM. The multiple panels of this figure highlight the influences coming from black hole electric charge, dark matter effects and nonlinear electrodynamics effects. (a) The upper left panel shows the critical shadow radius affected by PFDM in the absence of nonlinear electrodynamics effects (with $a=0$). (b) The upper right panel illustrates the black hole shadow in celestial coordinates affected by PFDM in the absence of nonlinear electrodynamics effects (with $a=0$ and $Q=0.5M$). (c) The middle left panel shows the critical shadow radius affected by PFDM in the presence of nonlinear electrodynamics effects (with $a=0.5M^{2}$). (d) The middle right panel illustrates the black hole shadow in celestial coordinates affected by PFDM in the presence of nonlinear electrodynamics effects (with $a=0.5M^{2}$ and $Q=0.5M$). (e) The lower left panel presents the critical shadow radius influenced by nonlinear electrodynamics effects (with $\lambda_{\text{DM}}=0.2M$). (f) The lower right panel shows the black hole shadow in celestial coordinates influenced by black hole electric charge (with $a=0.5M^{2}$ and $\lambda_{\text{DM}}=0.2M$). In all panels of this figure, the black hole shadow radius $r_{sh}$, black hole electric charge $Q$, PFDM parameter $\lambda_{\text{DM}}$ and nonlinear electrodynamics parameter $a$ are given in unit of black hole mass $M$.
@@
Interestingly, the black hole shadow results for Euler-Heisenberg black hole surrounded by PFDM in the presence and absence of nonlinear electrodynamics effects / QED effects could also exhibit a slight difference. In the absence of nonlinear electrodynamics effects, because of the emergence of naked singularity when black hole charge $Q>Q_{\text{extreme}}$, the black hole shadow disappears for large black hole electric charge. On the other hand, in the presence of nonlinear electrodynamics effects, although the Euler-Heisenberg black hole (surrounded by PFDM) does not produce any naked singularities (see Appendix B for the discussions on horizons), the black hole shadow may still disappear under some circumstances. In the middle left panel of figure 6, when PFDM parameter takes the values $\lambda_{\text{DM}}=0.0M,0.1M,0.2M$, the critical shadow radius calculated using expression (30) is not continuous when black hole charge reaches a critical value $Q=Q_{\text{discontinuity}}$. This is caused by the discontinuity of unstable photon sphere positions in the vicinity of Euler-Heisenberg black hole. When black hole charge exceeds this critical value, the number of unstable photon sphere around Euler-Heisenberg black hole undergoes a sudden change  [^3] When black hole charge satisfies $Q<Q_{\text{discontinuity}}$, there are two unstable photon spheres near the Euler-Heisenberg black hole (surrounded by PFDM), and the outer branch of unstable photon sphere determines the black hole shadow size via expression (30). When black hole charge becomes $Q>Q_{\text{discontinuity}}$, there is only one unstable photon sphere. The outer branch of unstable photon spheres in the $Q<Q_{\text{discontinuity}}$ cases is disappeared, which makes the unstable photon sphere position undergo a sudden change when black hole charge passes the critical point $Q=Q_{\text{discontinuity}}$.. Once the black hole charge comes across the discontinuity point (namely $Q>Q_{\text{discontinuity}}$), it is questionable that the results obtained using expression (30) could still represent a conventional “black hole shadow” or not. This issue deserves a more deeply study in the future.
5 Conclusions
@@
In this work, the gravitational lensing of Euler-Heisenberg black hole surrounded by PFDM is studied. This kind of black hole solution, which is solved from the gravitational field equation coupled with nonlinear electromagnetic field and dark matter field, could give rise to the nonlinear electrodynamics effects and PFDM effects in charge black hole systems. It provides us a simple way to investigate the interplay of nonlinear electrodynamics effects (or QED effects) and dark matter effects on the studies of black holes. These effects coming from nonlinear electrodynamics and dark matter can be parameterized by parameters $a$ and $\lambda_{\text{DM}}$ respectively. In the presented work, we mainly focus on the gravitational lensing observables for Euler-Heisenberg black hole surrounded by PFDM. The gravitational deflection angle of light, time delay of light, precession angle and trajectories of massive object’s bound orbits, black hole shadow are calculated numerically from geodesics.
@@
Our numerical results presented in this work show that influences coming from black hole electric charge, nonlinear electromagnetic effects, dark matter effects exhibit a similar tendency for gravitational lensing observables (such as gravitational deflection angle of light and time delay of light), precession angle of massive object’s bound orbits, and the black hole shadow size. Firstly, the Euler-Heisenberg black hole surround by PFDM with a larger electric charge could reduce the gravitational deflection angle, time delay of light, precession angle of bound orbits and black hole shadow radius. Secondly, the nonlinear electrodynamics effects parameterized by $a$ has very tiny influences on these quantities. Thirdly, the most significant impacts on these observables are coming from the dark matter effects. The gravitational deflection angle, time delay of light, precession angle of bound orbits and black hole shadow radius are all greatly changed when PFDM parameter $\lambda_{\text{DM}}$ is varied. Particularly, a larger PFDM parameter could greatly reduce the gravitational deflection angle, time delay of light and precession angle of bound orbits. This can be explained from the spacetime metric function (6) for Euler-Heisenberg black hole surrounded by PFDM, in which the PFDM parameter plays a role of “effective mass”. A positive PFDM parameter contributes to a negative “effective mass” in the spacetime metric, therefore it can diminish the gravitational lensing observables (such as gravitational deflection angle and time delay of light). Moreover, the dark matter also has notable influences on the shape and eccentricity of massive object’s bound orbits. For black hole shadow size, when PFDM parameter is smaller than a critical value, the black hole shadow radius gets reduced as the increasing of PFDM parameter. When PFDM parameter exceeds this critical value, the black hole shadow radius could be magnified when PFDM parameter increases. The critical value of PFDM parameter is roughly $0.5M<\lambda_{\text{critical}}<0.8M$ for different black hole charge values.

### v4

@@ 4.4 Black Hole Shadow
Figure 6: The black hole shadow of Euler-Heisenberg black hole surrounded by PFDM. (a) The upper left panel shows the critical shadow radius affected by PFDM and black hole charge in the absence of nonlinear electrodynamics effects. (b) The upper right panel illustrates the black hole shadow affected by PFDM in the absence of nonlinear electrodynamics effects. (c) The middle left panel shows the critical shadow radius affected by PFDM and black hole charge in the presence of nonlinear electrodynamics effects. (d) The middle right panel illustrates the black hole shadow affected by PFDM in the presence of nonlinear electrodynamics effects. (e) The lower left panel presents the critical shadow radius influenced by black hole charge and nonlinear electrodynamics effects. (f) The lower right panel shows the black hole shadow influenced by black hole electric charge. In all panels, the horizontal and vertical axes are given in unit of black hole mass.
@@
Interestingly, the black hole shadow results for Euler-Heisenberg black hole surrounded by PFDM in the presence and absence of nonlinear electrodynamics effects / QED effects could also exhibit a slight difference. In the absence of nonlinear electrodynamics effects, because of the emergence of naked singularity when black hole charge $Q>Q_{\text{extreme}}$, the black hole shadow disappears for large black hole electric charge. On the other hand, in the presence of nonlinear electrodynamics effects, the black hole shadow behavior becomes more complicated. When the nonlinear electrodynamics parameter takes a negative value ($a<0$), the Euler-Heisenberg black hole (surrounded by PFDM) still possesses naked singularity for a large black hole charge. In such cases, the shadow for Euler-Heisenberg black hole disappears when black hole charge exceeds its extreme value $Q>Q_{\text{extreme}}$ (with the extreme charge a little bit larger than the RN black hole case), which is shown in the lower left panel of figure 6. However, when the nonlinear electrodynamics parameter takes a positive value ($a>0$), the Euler-Heisenberg black hole (surrounded by PFDM) does not produce any naked singularities (see Appendix A for the discussions on horizons), the black hole shadow calculated from expression (30) can still appear for a larger black hole charge, but the critical shadow radius may undergo a discontinuity under some circumstances. In the middle left panel of figure 6, when PFDM parameter takes the values $\lambda_{\text{DM}}=0.0M,0.1M,0.2M$, the critical shadow radius calculated using expression (30) is not continuous when black hole charge reaches a critical value $Q=Q_{\text{discontinuity}}$. This is caused by the discontinuity of unstable photon sphere positions in the vicinity of Euler-Heisenberg black hole. When black hole charge exceeds this critical value, the number of unstable photon spheres around Euler-Heisenberg black hole undergoes a sudden change  [^4] When black hole charge satisfies $Q<Q_{\text{discontinuity}}$, there are two unstable photon spheres near the Euler-Heisenberg black hole (surrounded by PFDM), and the outer branch of unstable photon sphere determines the black hole shadow size via expression (30). When black hole charge becomes $Q>Q_{\text{discontinuity}}$, there is only one unstable photon sphere. The outer branch of unstable photon spheres in the $Q<Q_{\text{discontinuity}}$ cases is disappeared, which makes the unstable photon sphere position undergo a sudden change when black hole charge passes the critical point $Q=Q_{\text{discontinuity}}$.. Once the black hole charge comes across the discontinuity point (namely $Q>Q_{\text{discontinuity}}$), it is questionable that the results obtained using expression (30) could still represent a conventional “black hole shadow” or not. This issue deserves a more deeply study in the future.
5 Conclusions and Perspectives
@@
In this work, the gravitational lensing of Euler-Heisenberg black hole surrounded by PFDM is studied. This kind of black hole solution, which is solved from the gravitational field equation coupled with nonlinear electromagnetic field and dark matter field, could give rise to the nonlinear electrodynamics effects and PFDM effects in charge black hole systems. It provides us a simple way to investigate the interplay of nonlinear electrodynamics effects (or QED effects) and dark matter effects on the studies of black holes. These effects coming from nonlinear electrodynamics and dark matter can be parameterized by parameters $a$ and $\lambda_{\text{DM}}$ respectively. Particularly, the nonlinear electrodynamics effects have significant impact on the horizon structure of charged black hole. In the presented work, we mainly focus on the gravitational lensing observables for Euler-Heisenberg black hole surrounded by PFDM. The gravitational deflection angle of light, time delay of light, precession angle and trajectories of massive object’s bound orbits, black hole shadow are calculated numerically from geodesics.
@@
Our numerical results presented in this work show that influences coming from black hole electric charge, nonlinear electromagnetic effects, dark matter effects exhibit a similar tendency for gravitational lensing observables (such as gravitational deflection angle of light and time delay of light), precession angle of massive object’s bound orbits, and the black hole shadow size. Firstly, the Euler-Heisenberg black hole surround by PFDM with a larger electric charge could reduce the gravitational deflection angle, time delay of light, precession angle of bound orbits and black hole shadow radius. Secondly, the nonlinear electrodynamics effects parameterized by $a$ has very tiny influences on these quantities. Thirdly, the most significant impacts on these observables are coming from the dark matter effects. The gravitational deflection angle, time delay of light, precession angle of bound orbits and black hole shadow radius are all greatly changed when PFDM parameter $\lambda_{\text{DM}}$ is varied. Particularly, a larger PFDM parameter could greatly reduce the gravitational deflection angle, time delay of light and precession angle of bound orbits. This can be explained from the spacetime metric function (6) for Euler-Heisenberg black hole surrounded by PFDM, in which the PFDM parameter plays a role of “effective mass”. A positive PFDM parameter contributes to a negative “effective mass” in the spacetime metric, therefore it can diminish the gravitational lensing observables (such as gravitational deflection angle and time delay of light). Moreover, the dark matter also has notable influences on the shape and eccentricity of massive object’s bound orbits. For the black hole shadow size, when PFDM parameter is smaller than a critical value, the black hole shadow radius gets reduced as the increasing of PFDM parameter. When PFDM parameter exceeds this critical value, the black hole shadow radius could be magnified when PFDM parameter increases. The critical value of PFDM parameter is roughly $0.5M<\lambda_{\text{critical}}<0.8M$ for different black hole charge values.
@@
Furthermore, there are higher order QED corrections go beyond Euler-Heisenberg theory, which may contribute to the $1/r^{b}$ (with $b>6$) term in the spacetime metric of charged black hole systems. Black hole solutions with higher order QED contributions beyond the Euler-Heisenberg effective theory would be deduced and solved from a fundamental quantum gravity theory, and those black hole solutions may be more complicated than the Euler-Heisenberg black hole presented in this work. The seeking of such black hole solutions with the higher order QED corrections deserves future studies, which could provide us insights into the quantum effects of black hole systems and enhance our understanding of gravitational and electromagnetic interactions.
@@ Acknowledgment
Appendix A Horizons for Euler-Heisenberg Black Hole Surrounded by Perfect Fluid Dark Matter
In the absence of nonlinear electrodynamics effects / QED effects and dark matter effects (with $a\neq 0$ and $\lambda_{\text{DM}}=0$), the spacetime reduces to a conventional RN spacetime, resulting in two horizons for $Q<M$, a single horizon for the extreme case $Q=M$, and a naked singularity at the center for $Q>M$. However, in the presence of nonlinear electrodynamics effects and dark matter effects, the horizon structure can be strongly influenced. In figure 8, we plot the horizon radius for Euler-Heisenberg black hole surrounded by PFDM, which is calculated by numerically solving equation (31). The multiple panels of this figure highlight the influences coming from dark matter effects and nonlinear electrodynamics effects.
Figure 8: The horizons of Euler-Heisenberg black hole surrounded by PFDM. (a) The upper left panel shows the charged black hole’s horizons affected by PFDM parameter $\lambda_{\text{DM}}$ without the nonlinear electrodynamics effects. (b) The upper right panel shows the Euler-Heisenberg black hole’s horizons affected by nonlinear electrodynamics effects without PFDM. (c-d) The lower left and lower right panels plot the horizon radius of Euler-Heisenberg black hole in the presence of dark matter and nonlinear electrodynamics (with $a=0.2M^{2}$ and $a=0.8M^{2}$ respectively). In all panels, the horizontal and vertical axes are given in unit of black hole mass.
In the absence of nonlinear electrodynamics effects (with $a=0$), the dark matter effects do not have obvious influences on the structure of horizons, which can be easily observed in the upper left panel of figure 8. Without the nonlinear electrodynamics effects, charged black hole surrounded by PFDM still possesses two horizons (inner and outer horizons) when $0<Q<Q_{\text{extreme}}$. When the black hole charge exceeds the extremal black hole charge ($Q>Q_{\text{extreme}}$), the spacetime becomes horizonless and produces a naked singularity in the center $r=0$. This is similar to the naked singularity in conventional RN spacetime. In the presence of PFDM, the extremal black hole charge $Q_{\text{extreme}}$ becomes less than the conventional RN case. This can be explained from the “effective mass” role that PFDM played in spacetime metric (6), as we have explained in section 4. A positive PFDM parameter contributes to a negative “effective mass” in spacetime metric, and the corresponding extremal black hole charge is less than black hole mass ($Q_{\text{extreme}}<M$) for a given positive PFDM parameter ($\lambda_{\text{DM}}>0$).
In the presence nonlinear electrodynamics effects (with $a\neq 0$), the horizon structure for charged Euler-Heisenberg black holes can be dramatically changed, which is clearly shown in the upper right panel of figure 8. When the nonlinear electrodynamics parameter take a negative value $a<0$, the horizons for Euler-Heisenberg black hole are similar to the RN black hole, possessing inner and outer horizons for black hole charge $Q<Q_{\text{extreme}}$ and giving rise to naked singularity when $Q>Q_{\text{extreme}}$. However, for a positive nonlinear electrodynamics parameter ($a>0$), the horizons for charged black hole have three main differences compared to the RN black hole. Firstly, the naked singularity no longer exists when nonlinear electrodynamics parameter takes a positive value ($a>0$). In such cases, there is at least one horizon in Euler-Heisenberg black hole for any given black hole charge. Secondly, for a small black hole charge, the Euler-Heisenberg black hole has only one horizon, unlike the two horizon behavior (the inner and outer horizons) for pure RN spacetime with $a=0$. In the upper right panel of figure 8, various curves correspond to positive nonlinear electrodynamics parameter values ($a>0$) in the branch $M<r_{H}<2M$ and $Q<M$ are almost overlapped with each other, but the RN black hole with $a=0$ has a second branch at $0<r_{H}<M$ and $Q<M$ (labeled by black solid curve in the upper right panel). Thirdly, if the black hole electric charge falls into the region $0.85M<Q<1.02M$, the Euler-Heisenberg black hole with a small positive nonlinear electrodynamics parameter ($0<a<0.5M^{2}$) could have three horizons. However, a larger nonlinear electrodynamics parameter (for instance, the $a=0.65M^{2}$ and $a=0.8M^{2}$ in the upper right panel of figure 8) could prevent the occurrence of three horizons in Euler-Heisenberg black hole, producing a single horizon for arbitrary electric charge values.
When taking into account both nonlinear electrodynamics effects (with a positive nonlinear electrodynamics parameter $a>0$) and dark matter effects (with $\lambda_{\text{DM}}>0$), the horizons of Euler-Heisenberg black hole surrounded by PFDM are plotted in the lower panels of figure 8, for nonlinear electrodynamics parameters $a=0.2M^{2}$ and $a=0.8M^{2}$ respectively  [^5] When the nonlinear electrodynamics parameter becomes negative, the horizon behavior for Euler-Heisenberg black hole surrounded by PFDM is similar to the RN black hole surrounded by PFDM given in the upper left panel of figure 8. Comparing the positions of horizon for Euler-Heisenberg black hole without dark matter (in the upper right panel) and those in the presence of dark matter (in the lower left and lower right panels), we can obtain the following conclusions. In the presence of dark matter and nonlinear electrodynamics, the naked singularity is no longer emerged for arbitrary black hole electric charge, which is the same as the results for Euler-Heisenberg black hole without PFDM. The dark matter influences the horizon size and changes the width of three-horizon region in the parameter space of black hole charge. A positive PFDM parameter ($\lambda_{\text{DM}}>0$) results in a smaller horizon radius, which can be explained using the “effective mass” role of the PFDM played in spacetime metric. Furthermore, a positive PFDM parameter narrows the three-horizon region in the parameter space of black hole charge. For instance, when nonlinear electrodynamics parameter gets the value $a=0.2M^{2}$, the three-horizon region becomes narrowed when $0<\lambda_{\text{DM}}\leq 0.3M$ (compared with those without PFDM plotted in the upper right panel), while the $\lambda_{\text{DM}}=0.5M$ case does not possess a three-horizon region. For a large nonlinear electrodynamics parameter (such as $a=0.8M^{2}$ in the lower right panel), the Euler-Heisenberg black hole in the presence of dark matter does not possess a three-horizon region, no matter what the PFDM parameter is varied.